OpenStudy (anonymous):

cos^-1(cos4pi/5) can someone help me with this one? And explain how they got the answer as well Will fan and medal

OpenStudy (tkhunny):

Cosine and Inverse Cosine are inverses. Each un-does what the other did in the first place. The obvious answer should be 4pi/5, but there could be more to it. What do you think?

OpenStudy (tkhunny):

Rephrase: IF you see that they are inverses, the result should be obvious. It's a big if. Do you see it?

OpenStudy (anonymous):

sin^-1(sin(-pi/10) @tkhunny Would the answer be -pi/10 then?

OpenStudy (anonymous):

may i help

OpenStudy (anonymous):

sure

OpenStudy (tkhunny):

Maybe. The technical problem is the Inverse Sine and Cosine Functions is their range. Whatever the structure, it has to give you something in this range. You might thing it's just -pi/10, but that is it? In this case, \(\sin^{−1}(\sin(−π/10))=-π/10\), an angle in the right range but having the same sine. Interestingly, \(\cos^{−1}(\cos(−π/10))=+π/10\)